# Choosing alpha post hoc: the danger of multiple standard significance thresholds

### v1

@@ 1.1 The call to decrease significance thresholds
Recently, there has been an intense debate in many scientific fields on whether the significance threshold for declaring scientific discoveries should be changed from $0.05$ to $0.005$ (Borenstein, 2001; Ioannidis, 2005; Banerjee et al., 2009; Mudge et al., 2012; Martin and Bridgmon, 2012; Kim and Ji, 2015; Ioannidis, 2016; Szucs and Ioannidis, 2017; Benjamin et al., 2018; Ioannidis, 2018; Held, 2019; Di Leo and Sardanelli, 2020; Khan et al., 2020; Aguinis et al., 2021; Schnog et al., 2021; Thakur and Jha, 2022; Maier and Lakens, 2022; Fitzpatrick et al., 2024). In most statistics textbooks, the “significance threshold” or “significance level” is denoted by $\alpha$ and is assumed to be prespecified, i.e., data-independent (Van der Vaart, 1998; Lehmann and Romano, 2022; Hansen, 2022). In particular, $\alpha$ is assumed independent of the p-value, since otherwise $\alpha$ would depend on the data (Hubbard and Bayarri, 2003). One important paper that stimulated the mentioned debate, bears the title “Redefine statistical significance” (Benjamin et al., 2018). In this article, 72 influential scientists propose changing the “threshold for statistical significance for claims of new discoveries” to $0.005$.
As a consequence of this proposal, scientists may find themselves in a field where there is not one standard threshold, but multiple: some groups of scientists may support $\alpha=0.005$ and some $\alpha=0.05$. This can also translate into the implicit or explicit preferences of journals, with some journals preferring $0.005$ and others $0.05$.
@@ 1.2 The statistical danger of differing thresholds
The message of this paper is that if within one field, there are multiple standard significance thresholds, then researchers may be tempted to first analyze their data and then choose the $\alpha$ that suits them. The problem can best be indicated with an example.
Suppose a researcher is interested in a certain null hypothesis and plans to compute a p-value for it. In her field there are two standard significance thresholds, $0.005$ and $0.05$, simply because some researchers or journal editors in her field advocate 0.005 and others advocate 0.05. After she has computed the p-value, she wants to share it in some form, e.g. by presenting it at a conference or submitting it to a journal. Then the choice of where to present or where to submit her results will potentially depend on the p-value that she finds. For example, if she finds $p=0.02$, then she may be more likely to submit to a journal that accepts $\alpha=0.05$ than to a journal that prefers $\alpha=0.005$. On the other hand, if she finds an extremely small p-value, then she may prefer a journal with $\alpha=0.005$, since then her finding is significant at a more stringent threshold. Consequently, it would be tempting for the researcher not to fix her $\alpha$ before running her experiment, but to let her $\alpha$ depend on the journal or conference, which she chooses after obtaining her p-value.
The statistical issue that arises, is that the researcher’s $\alpha$ is no longer independent of her p-value. Consequently, the test loses its usual decision-theoretic interpretation, i.e., there is no guarantee anymore that the type I error probability is below $\alpha$ (Hubbard and Bayarri, 2003; Grünwald, 2024; Koning, 2024a). Note that this issue is about more than publication bias. It is about the validity of an individual test, from the perspective of the individual researcher: when she lets $\alpha$ depend on the data, she cannot validly interpret her own statistical test anymore.
In theory, a solution would be that the researcher fixes her significance threshold beforehand and never changes it. However, one can easily imagine a researcher who finds $p=0.003$ and then claims her significance level is $\alpha=0.005$, even though she never fixed $\alpha$ beforehand.
@@ 1.3 A related discussion: abandoning significance thresholds
In parallel to the discussion on what the threshold should be, there has been a discussion on whether significance thresholds should be used at all (Hubbard and Lindsay, 2008; Wasserstein and Lazar, 2016; McShane et al., 2019; Amrhein et al., 2019). Indeed, a downside of significance testing in practice is that it may overemphasize a specific p-value threshold rather than interpreting p-values continuously, considering effect sizes, and using other scientifically relevant information (Borenstein, 1994; Greenland et al., 2016; Thompson, 2007; Betensky, 2019; Scheel et al., 2021; Imbens, 2021; Muff et al., 2022). For example, the famous ASA statement on p-values (Wasserstein and Lazar, 2016) notes that “Scientific conclusions and business or policy decisions should not be based only on whether a p-value passes a specific threshold.” Note that the ASA statement is not against p-values, but against too much focus on p-value thresholds.
Many statisticians agree with the ASA statement’s point, but nevertheless we see that in the recent statistical literature, significance testing is still prominent (Lehmann and Romano, 2022; Hansen, 2022). One reason is that from a mathematical point of view, there is often essentially no difference between a formula for a significance test, a formula for a p-value and a formula for a confidence interval. For example, confidence intervals can often be directly derived by ‘inverting’ a test. Another reason is that some statisticians still value type I error control, which is only possible if some $\alpha$ is chosen. Hence, although many statisticians largely agree with the ASA statement, they continue to develop hypothesis tests.
Beyond theoretical statistics, in science in general we see that significance testing has not disappeared (Hubbard, 2019; Mayo and Hand, 2022; Thakur and Jha, 2022; Maier and Lakens, 2022; Hansen, 2022; Fitzpatrick et al., 2024). Moreover, if one abandons significance tests, then one is still likely to compute confidence intervals, which also require choosing $\alpha$. In that case, choosing $\alpha$ post hoc causes issues similar to those discussed in this paper.
@@ 1.4 This paper
Throughout the paper, we stay within the classical Neyman-Pearson theory of hypothesis testing. We note that one of the main reasons why classical Neyman-Pearson hypothesis testing works in practice, is that there exists a single standard level $\alpha$ in most fields. This automatically avoids that $\alpha$ is chosen based on the data. Although it is known that having $\alpha$ depend on the p-value is invalid (Hubbard and Bayarri, 2003), this paper seems to be the first to point out this will likely occur when within a field there are multiple standard thresholds.
@@
Further, we discuss in detail what can go wrong in terms of type I error control. We evaluate what the type I error rate is, conditional on a certain $\alpha$ being chosen based on the p-value. Apart from discussing the situation where the researcher chooses between two significance levels, we will also consider the most extreme scenario where there are many choices for $\alpha$.
Moreover, for a test with a data-dependent level $\alpha$, we consider a measure of evidence for how problematic this is on average. If this measure is large, this means the conditional type I error rate of the test is much too large in an average sense. In the examples that we provide in Sections 3 and 4, the interpretation of the hypothesis test is ruined due to the post hoc selection of $\alpha$.
Throughout, we focus on the case where one hypothesis is tested. In case there are multiple hypotheses and a multiple testing correction is used, similar issues arise when $\alpha$ is chosen post hoc. This is discussed in the Supplementary Material. In Section 5 and the Discussion, we briefly point out some potential solutions to the discussed issue.
2 Notation and concepts
2.1 Tests with data-dependent $\alpha$
Consider any hypothesis $H_{0}$ and a corresponding hypothesis test $\phi_{\alpha}$. The subscript is used because the test depends on the chosen significance level $\alpha>0$. The test $\phi_{\alpha}$ is a function from the sample space to the set $\{0,1\}$, where $1$ indicates that $H_{0}$ is rejected and 00 indicates that $H_{0}$ is not rejected. We will indicate the data with $X$, which takes values in the sample space. We view the test $\phi_{\alpha}=\phi_{\alpha}(X)$ as a function of $X$, i.e., given $\alpha$, the outcome of the test $\phi_{\alpha}$ is entirely determined by the data $X$.
In classical hypothesis testing, the level $\alpha$ is always assumed to be selected independently from the data. For such an independently selected $\alpha$, it is typical to use a test $\phi_{\alpha}$ with the guarantee that it does not reject the hypothesis too often when it is true, namely with probability at most $\alpha$. We assume throughout that if $\alpha$ is chosen independently from $X$, then the tests we consider are exactly valid, in the sense that $\mathbb{P}(\phi_{\alpha}=1)=\alpha$ if $H_{0}$ is true.
As indicated in the introduction, there are scenarios where $\alpha$ would depend on $X$ – for example if $\alpha$ depends on the p-value. Then $\alpha$ is a function of $X$, although we will often write $\alpha$ instead on $\alpha(X)$ for short. If $\alpha$ is data-dependent, then conditional on $\alpha$ taking a certain value, we possibly lose the guarantee that $\mathbb{P}_{H_{0}}(\phi_{\alpha}=1)\leq\alpha$. For example, suppose $\alpha$ takes two possible values, e.g. 0.005 and 0.05, depending on the data $X$. Then, depending on how $\alpha$ exactly depends on $X$, it can be the case that $\mathbb{P}({\phi_{\alpha}=1|\alpha=0.005})$ is not at most $0.005$, but larger than $0.005$. This will be illustrated in Section 3.
2.2 Measuring the discrepancy between $\alpha$ and the type I error rate
@@
Let $A\subseteq[0,1]$ denote the set of possible values that $\alpha(X)$ can take. Throughout we use ‘$a$’ to denote a value from $A$, independent of the data. We can then ask whether conditional on the data-dependent $\alpha$ taking the value $a$, the type I error rate is below $\alpha$, i.e. whether $\mathbb{P}_{H_{0}}(\phi_{\alpha}=1|\alpha=a)\leq a$. A measure of the discrepancy between $a$ and $\mathbb{P}_{H_{0}}(\phi_{\alpha}=1|\alpha=a)$ is the difference
$d_{a}:=\mathbb{P}\big{(}\phi_{\alpha}=1|\alpha=a\big{)}-a.$ (1)
@@
Ideally, we would have $d_{a}=0$ for all $a\in A$, but this is often not the case for many $a\in A$, if $\alpha$ depends on the data. This is well known (Hubbard and Bayarri, 2003) and illustrated in Sections 3 and 4. Now suppose that for some $a\in A$, the difference is $d_{a}=0.01$. Do we consider this difference small or large? To a large extent, this depends on what $a$ is. For example, if $a=0.05$, then having $d_{a}=0.01$ is not highly problematic, since $\mathbb{P}(\phi_{\alpha}=1|\alpha=a)=0.06$ is then not much larger than $a$, in relative terms. However, if $a=0.005$, then $d_{a}=0.01$ implies that the true error rate $\mathbb{P}_{H_{0}}(\phi_{\alpha}=1|\alpha=a)=0.015$ is three times larger than $a$. In that case, $d_{a}=0.01$ is a huge difference. Thus, $d_{a}$ is arguably not the best measure of how liberal the test is. Instead it can be more meaningful to look at the relative discrepancy
$r_{a}:=\mathbb{P}_{H_{0}}\big{(}\phi_{\alpha}=1|\alpha=a\big{)}/a,$ (2)
@@
where we remind the reader that $\alpha=\alpha(X)$ depends on the data. We will call $r_{a}$ the discrepancy ratio conditional on $\alpha=a$. We would like this ratio to be at most 1. If $r_{a}<1$, then the test is conservative conditional on $\alpha=a$, and if $r_{a}>1$ the test is liberal conditional on $\alpha=a$. Note that for a given number $a$, $r_{a}$ is a fixed number. In contrast, since $\alpha=\alpha(X)$ depends on the data, $r_{\alpha}$ is data-dependent.
2.3 Expected discrepancy
If the discrepancy ratio $r_{a}$ is below 1 for every $a\in A$, then in particular the expected value of the discrepancy ratio will be below 1, i.e.,
$\mathbb{E}r_{\alpha}=\mathbb{E}\big{\{}\mathbb{P}(\phi_{\alpha}=1|\alpha)/\alpha\big{\}}$ (3)
will be at most 1. Note that this is the expected value of a conditional probability divided by $\alpha$. Likewise, if (3) is larger than 1, it follows that there is at least one $a\in A$ for which (2) is larger than 1.
@@ 2.3 Expected discrepancy
Note that in general, if $\mathbb{E}r_{\alpha}$ is much larger than $1$ under $H_{0}$, it follows that either there is at least a small probability that $r_{\alpha}$ is much larger than $1$, or there is a large probability that $r_{\alpha}$ is at least somewhat larger than 1. Both these possibilities are problematic for researchers. Thus, $\mathbb{E}r_{\alpha}$ is a summary measure of how wrong things go on average. If $\mathbb{E}r_{\alpha}\gg 1$, this indicates that the test procedure is seriously flawed, in an average sense.
@@
It is useful to note that $\mathbb{E}r_{\alpha}$ can be rewritten in the following ways:
$\mathbb{E}r_{\alpha}=\mathbb{E}\big{\{}\mathbb{P}(\phi_{\alpha}=1|\alpha)/\alpha\big{\}}=\mathbb{E}\big{\{}\mathbb{E}\big{(}\phi_{\alpha}|\alpha\big{)}/\alpha\big{\}}=$
$\mathbb{E}\big{\{}\mathbb{E}\big{(}\phi_{\alpha}/\alpha|\alpha\big{)}\big{\}}=\mathbb{E}\big{(}\phi_{\alpha}/\alpha\big{)}.$ (4)
The equalities (4) tell us that $\mathbb{E}r_{\alpha}$ can be rewritten as the expectation of a conditional expectation, which can in turn be rewritten as a simple expectation.
3 Example 1: two possible significance levels $\alpha$
Consider a researcher who tests one hypothesis. In her field there are two standard significance thresholds, $a_{1}$ and $a_{2}$, where $10<a_{1}<a_{2}<$ 1. As explained in the Introduction, there are scenarios where her $\alpha$ would depend on the p-value. In particular, if the researcher obtains a p-value $p$ with $a_{1}<p\leq a_{2}$, she may be more likely to take $\alpha=a_{2}$ than when $p\leq a_{1}$, in which case she could be more likely to take $\alpha=a_{2}$. As a simple behavioural model, suppose that the researcher takes $\alpha=a_{1}$ if $p\leq a_{1}$ and takes $\alpha=a_{2}$ if $a_{1}<p\leq 1$.
We will use the notation from Section 2.1. Since the threshold depends on the p-value, it depends on the data. Thus, $\alpha(X)$ depends on $X$, although we supress ‘$(X)$’ for brevity. We have $\phi_{\alpha}=\phi_{\alpha(X)}=1$ if and only if $p\leq\alpha$. If $\alpha$ would be fixed, then the test $\phi_{\alpha}$ would be valid in the sense that $\mathbb{P}(\phi_{\alpha}=1)\leq\alpha$. However, because $\alpha$ depends on the p-value, the interpretation of the test is ruined, as we will see.
3.1 No type I error control conditional on $\alpha$ or marginally
@@
We investigate the type I error probability conditional on a certain value for $\alpha$ being chosen. What happens conditional on $\alpha=a_{1}$ can be seen immediately: if $\alpha=a_{1}$, then this means that $p$ was apparently at most $a_{1}$. This means that $H_{0}$ is rejected. Thus, under $H_{0}$, conditional on $\alpha=a_{1}$, the type I error probability is not $a_{1}$, but 1. Further, the type I error probability is larger than $a_{1}$ marginally. Thus, the familiar interpretation of the hypotheses test is ruined. This means that the researcher cannot validly interpret her own statistical test anymore, in the usual decision-theoretic way.
The following proposition states what happens to the type I error probability conditional on $\alpha=a_{2}$. The proof is in the appendix.
Proposition 1.
For the above procedure that takes $\alpha=a_{1}$ if $p\leq a_{1}$ and $\alpha=a_{2}$ otherwise, under $H_{0}$ we have
$\mathbb{P}(\phi_{\alpha}=1|\alpha=a_{2})=\big{(}(1-a_{1}/a_{2})/(1-a_{1})\big{)}a_{2}<a_{2}.$
Thus, conditional on the data-dependent $\alpha$ being $a_{2}$, the type I error rate does not exceed $a_{2}$, although the test is conservative since $\mathbb{P}(\phi_{\alpha}=1)$ is strictly smaller than $a_{2}$. For example, for $a_{1}=0.005$ and $a_{2}=0.05$, we have $\mathbb{P}(\phi_{\alpha}=1|\alpha=a_{2})=a_{2}((1-0.005/0.05)/(1-0.005))\approx a_{2}\cdot 0.9045$. Thus, in the notation of Section 2.2, $r_{a_{2}}\approx 0.9045$.
In summary, conditional on $\alpha=a_{1}$ the type I error probability is too large and conditional on $\alpha=a_{2}$ the type I error probability is on the conservative side. Put differently, the discrepancy ratio is $r_{\alpha}>1$ if $\alpha=a_{1}$ and $r_{\alpha}<1$ if $\alpha=a_{2}$.
3.2 No type I error control in terms of expected discrepancy
@@
Next, we can ask whether perhaps we control the type I error probability in expectation, i.e., whether $\mathbb{E}r_{\alpha}\leq 1$ under $H_{0}$. This would mean that at least in an average sense, we have type I error control. The following proposition implies that this does not hold.
Proposition 2.
For the procedure that takes $\alpha=a_{1}$ if $p\leq a_{1}$ and $a_{2}$ otherwise, under $H_{0}$ we have
$\mathbb{E}r_{\alpha}=1+(a_{2}-a_{1})/a_{2}>1.$
@@
For example, if $a_{1}=0.005$ and $a_{2}=0.05$, then $\mathbb{E}_{H_{0}}(r_{\alpha})=1.9$. Since $\mathbb{E}_{H_{0}}(r_{\alpha})$ is not bounded by 1, it follows that $r_{\alpha}$ is larger than 1 conditional on at least one possible value of $\alpha$. We already saw that this is the case for $\alpha=a_{1}$. We conclude that both conditionally and on average, the researcher’s test does not provide type I error control.
4 Example 2: Infinitely many significance levels $\alpha$

### v2

@@ 1.1 The call to decrease significance thresholds
Conducting a hypothesis test requires choosing a “significance level” or “significance threshold”, typically denoted by $\alpha$ [48, 32, 13]. It is assumed to be prespecified, and therefore independent of the p-value [17]. Recently, there has been an intense debate in many fields on whether the threshold should be changed from 0.05 to 0.005 [7, 20, 3, 37, 34, 26, 21, 43, 4, 22, 29, 14, 8, 25, 1, 41, 44, 9]. In one important paper that stimulated the debate, 72 influential scientists propose changing the “threshold for statistical significance for claims of new discoveries” to 0.005 [4].
As a consequence of such proposals, scientists may find themselves in a field where there is not one standard threshold, but multiple: some groups of scientists may support $\alpha=0.005$ and some $\alpha=0.05$. This can also translate into the implicit or explicit preferences of journals, with some journals preferring $0.005$ and others $0.05$.
@@ 1.2 The statistical danger of differing thresholds
The message of this paper is that if within one field, there are multiple standard significance thresholds, then researchers may be tempted to first analyze their data and then choose the threshold $\alpha$ that suits them.
For example, suppose a researcher is interested in a certain null hypothesis and plans to compute a p-value for it. In her field there are two standard significance thresholds, $0.005$ and $0.05$, because some researchers or journal editors in her field advocate 0.005 and others advocate 0.05. After she has computed the p-value, she wants to share it in some form, e.g. by presenting it at a conference or submitting it to a journal. Then the choice of where to present or where to submit her results will potentially depend on the p-value that she finds. For example, if she finds $p=0.003$, then she may prefer a journal that supports $\alpha=0.005$ over a journal that supports $\alpha=0.05$, since then her finding is significant at a more stringent threshold. Consequently, it would be tempting for the researcher not to fix her $\alpha$ before running her experiment, but to let her $\alpha$ depend on the journal or conference, which she chooses after obtaining her p-value.
The statistical issue that arises, is that the researcher’s $\alpha$ is no longer independent of her p-value. Consequently, the test loses its usual decision-theoretic interpretation, i.e., there is no guarantee anymore that the type I error probability is below $\alpha$ [17, 12, 28]. Note that this issue is about more than publication bias. It is about the validity of an individual test, from the perspective of the individual researcher: when she lets $\alpha$ depend on the data, she cannot validly interpret her own statistical test anymore. In theory, a solution would be that the researcher fixes her significance threshold beforehand, but one can easily imagine a researcher who finds $p=0.003$ and then claims her significance level is $\alpha=0.005$, even though she never fixed $\alpha$ beforehand.
@@ 1.3 A related discussion: abandoning significance thresholds
In parallel to the discussion on what the threshold should be, there has been a discussion on whether significance thresholds should be used at all [18, 50, 36, 2]. Indeed, a downside of significance testing in practice is that it may overemphasize a specific p-value threshold rather than interpreting p-values continuously, considering effect sizes, and using other scientifically relevant information [6, 10, 45, 5, 40, 19, 38]. For example, the famous ASA statement on p-values [50] notes that “Scientific conclusions and business or policy decisions should not be based only on whether a p-value passes a specific threshold.” Note that the ASA statement is not against p-values, but against too much focus on p-value thresholds.
Many statisticians agree with the ASA statement’s point, but nevertheless we see that in the recent statistical literature, significance testing is still prominent [32, 13]. One reason is that from a mathematical point of view, there is often essentially no difference between a formula for a significance test, a formula for a p-value and a formula for a confidence interval. For example, confidence intervals can often be directly derived by ‘inverting’ a test. Another reason is that some statisticians still value type I error control, which is only possible if some $\alpha$ is chosen.
Beyond theoretical statistics, in science in general we see that significance testing has not disappeared [16, 35, 44, 33, 13, 9]. Moreover, if one abandons significance tests, then one is still likely to compute confidence intervals, which also require choosing $\alpha$.
@@ 1.4 This paper
Throughout the paper, we stay within the classical Neyman-Pearson theory of hypothesis testing. Our main point is that tests are often invalidated if $\alpha$ depends on the p-value [17], and that this will likely occur when within a field there are multiple standard thresholds. [46] make the following related point: one possible reason why a single $\alpha$ has been the norm for so long, is that it prevents researchers from choosing $\alpha$ post hoc.
@@
Further, we discuss in detail what can go wrong in terms of type I error control. We evaluate what the type I error rate is, conditional on a certain $\alpha$ being chosen based on the p-value. Moreover, we introduce a measure of how problematic a test with data-dependent $\alpha$ is. If this measure is large, this means the conditional type I error rate of the test is too large in an average sense. Apart from discussing the situation where the researcher chooses between two significance levels, we will also consider the most extreme scenario where there are many choices for $\alpha$. In Section 3 and the Discussion, we briefly point out some potential solutions to the discussed issue.
2 Discrepancies between $\alpha$ and the type I error rate
2.1 Notation and concepts
Consider a null hypothesis $H_{0}$ and a corresponding test $\phi_{\alpha}$, where $0<\alpha\leq 1$ indicates the chosen significance level. The test $\phi_{\alpha}=\phi_{\alpha}(X)$ is a function of the data $X$ and maps to the set $\{0,1\}$, where $1$ indicates that $H_{0}$ is rejected. In this paper, all mentioned probabilities and expectations will be under $H_{0}$.
@@ 2.1 Notation and concepts
Suppose that $\alpha$ possibly depends on $X$. Let $A\subseteq[0,1]$ denote the set of possible values that $\alpha$ can take. Throughout we use ‘$a$’ to denote a value from $A$, independent of the data. We can then ask whether conditional on the data-dependent $\alpha$ taking the value $a$, the type I error rate is below $\alpha$, i.e. whether $\mathbb{P}(\phi_{\alpha}=1\mid\alpha=a)\leq a$. A measure of the discrepancy between $a$ and $\mathbb{P}(\phi_{\alpha}=1\mid\alpha=a)$ is the difference
$d_{a}:=\mathbb{P}\big{(}\phi_{\alpha}=1\mid\alpha=a\big{)}-a.$
Ideally, we would have $d_{a}=0$ for all $a\in A$, but this is often not the case for many $a\in A$, if $\alpha$ depends on the data. This is well known [17] and illustrated in Sections 2.2 and 2.3.
@@
Now suppose that for some $a\in A$, the difference is $d_{a}=0.01$. Do we consider this difference small or large? To a large extent, this depends on what $a$ is. For example, if $a=0.05$, then having $d_{a}=0.01$ is not highly problematic, since $\mathbb{P}(\phi_{\alpha}=1\mid\alpha=a)=0.06$ is then not much larger than $a$, in relative terms. However, if $a=0.005$, then $d_{a}=0.01$ implies that the true error rate $\mathbb{P}(\phi_{\alpha}=1\mid\alpha=a)=0.015$ is three times larger than $a$. In that case, $d_{a}=0.01$ is a huge difference. Thus, $d_{a}$ is arguably not the best measure of how liberal the test is.
Instead it can be more meaningful to look at the relative discrepancy
$r_{a}:=\mathbb{P}\big{(}\phi_{\alpha}=1\mid\alpha=a\big{)}/a.$
@@
We will call $r_{a}$ the discrepancy ratio conditional on $\alpha=a$. We would like this ratio to be at most 1. If $r_{a}>1$, the test is liberal conditional on $\alpha=a$. Note that for a given number $a$, $r_{a}$ is a fixed number. In contrast, since $\alpha$ depends on the data, $r_{\alpha}$ is data-dependent.
@@
If the discrepancy ratio $r_{a}$ is below 1 for every $a\in A$, then in particular $\mathbb{E}r_{\alpha}$ will be below 1. If on the other hand $\mathbb{E}r_{\alpha}$ is much larger than $1$ under $H_{0}$, it follows that either there is at least a small probability that $r_{\alpha}$ is much larger than $1$, or there is a large probability that $r_{\alpha}$ is at least somewhat larger than 1. Both these possibilities are problematic for researchers. Thus, $\mathbb{E}r_{\alpha}$ is a summary measure of how wrong things go on average. If $\mathbb{E}r_{\alpha}\gg 1$, this indicates that the test procedure is seriously flawed, in an average sense.
@@
It is useful to note that $\mathbb{E}r_{\alpha}$ can be rewritten as a simple expectation:
$\mathbb{E}r_{\alpha}$ $=\mathbb{E}\big{\{}\mathbb{P}(\phi_{\alpha}=1\mid\alpha)/\alpha\big{\}}=\mathbb{E}\big{\{}\mathbb{E}\big{(}\phi_{\alpha}|\alpha\big{)}/\alpha\big{\}}$
$=\mathbb{E}\big{\{}\mathbb{E}\big{(}\phi_{\alpha}/\alpha\mid\alpha\big{)}\big{\}}=\mathbb{E}\big{(}\phi_{\alpha}/\alpha\big{)}.$ (1)
2.2 Example 1: two possible significance levels $\alpha$
Consider a researcher who tests one hypothesis. She computes a p-value, which we assume to exactly satisfy
$\mathbb{P}(p\leq\alpha)=\alpha$
under $H_{0}$ for every data-independent $\alpha$. Suppose that in her field there are two standard significance thresholds, $a_{1}$ and $a_{2}$, where $10<a_{1}<a_{2}<$ 1. As a simple behavioral model, assume she chooses the smaller threshold $\alpha=a_{1}$ if $p\leq a_{1}$ and chooses $\alpha=a_{2}$ if $a_{1}<p\leq 1$, so that $\alpha$ depends on $p$ and hence on the data.
@@ 2.2 Example 1: two possible significance levels $\alpha$
What happens to the type I error probability conditional on $\alpha=a_{1}$ can be seen immediately: if $\alpha=a_{1}$, then this means that $p$ was apparently at most $a_{1}$, which means that $H_{0}$ is rejected. Thus, under $H_{0}$, conditional on $\alpha=a_{1}$, the type I error probability is not $a_{1}$ but 1. Conditional on $\alpha=a_{2}$, the type I error probability is strictly smaller than $a_{2}$,
$\mathbb{P}(p\leq\alpha\mid\alpha=a_{2})$ $=\mathbb{P}(p\leq a_{2}\mid p>a_{1})$
$=\frac{a_{2}-a_{1}}{1-a_{1}}<\frac{a_{2}-a_{2}a_{1}}{1-a_{1}}=a_{2}.$
Thus, the discrepancy ratio is $r_{\alpha}>1$ if $\alpha=a_{1}$ and $r_{\alpha}<1$ if $\alpha=a_{2}$.
@@
Next, we can ask whether perhaps we control the type I error probability in expectation, in the sense that $\mathbb{E}r_{\alpha}\leq 1$ under $H_{0}$. This turns out to not be the case either, since under $H_{0}$, by (1),
$\mathbb{E}r_{\alpha}$ $=\mathbb{E}(\phi_{\alpha}/\alpha)$
$=\mathbb{P}(p\leq a_{1})\mathbb{E}(\phi_{\alpha}/\alpha\mid p\leq a_{1})$
$+\mathbb{P}(a_{1}<p\leq a_{2})\mathbb{E}(\phi_{\alpha}/\alpha\mid a_{1}<p\leq a_{2})$
$=a_{1}(1/a_{1})+(a_{2}-a_{1})(1/a_{2})$
$=1+(a_{2}-a_{1})/a_{2}>1.$
@@
For example, if $a_{1}=0.005$ and $a_{2}=0.05$, then $\mathbb{E}r_{\alpha}=1.9$. We conclude that both conditionally and on average, the researcher’s test does not provide type I error control.
2.3 Example 2: Infinitely many significance levels $\alpha$